OpenStudy (anonymous):

Solve the following system using the substitution method. x + 3y = -5 5x + 6y = -19

OpenStudy (campbell_st):

rewrite the 1st equation so that 3y = -5 - x and substitute it into the 2nd equation 5x + 2(3y) = -19 or 5x + 2(-5 - x )= -19 which becomes 5x - 10 - 2x = -19 simplifying gives 3x - 10 = -19 now solve for x. when you have a value for x just substitute into either equation to find y

OpenStudy (anonymous):

ok